OpenStudy (anonymous):

A ball is thrown upward from the top of a building. The function below shows the height of the ball in relation to sea level, ft, in feet, at different times t in seconds: f(t) = -16t2 + 32t + 90 The average rate of change of f(t) from t = 4 seconds to t = 6 seconds is _____ feet per second.

OpenStudy (anonymous):

@Jamierox4ev3r

Vocaloid (vocaloid):

we'll be using the formula for average rate of change [f(6)-f(4)] / (6-4) so, first start by finding f(6)-f(4) = ?

OpenStudy (anonymous):

6-4=2 @Vocaloid

Vocaloid (vocaloid):

not quite f(t) = -16t^2 + 32t + 90 f(6) = ?

OpenStudy (anonymous):

oh!! 90? @Vocaloid

Vocaloid (vocaloid):

almost, but not quite... t = 6 -16t^2 = ?

Vocaloid (vocaloid):

^ means exponent

OpenStudy (anonymous):

whoops i wrote it down wrong on my paper hold on @Vocaloid

OpenStudy (anonymous):

i got 858? @Vocaloid

OpenStudy (anonymous):

forgot the negative ugh hold on

OpenStudy (anonymous):

-294**** @Vocaloid

Vocaloid (vocaloid):

good, so f(6) = -294 now find f(4)

OpenStudy (anonymous):

how do i do that?

OpenStudy (anonymous):

@Vocaloid

Vocaloid (vocaloid):

using the same method you just did, just set t = 4 instead of 6

OpenStudy (anonymous):

okay hold on

OpenStudy (anonymous):

-38 @Vocaloid

Vocaloid (vocaloid):

good, now plug everything back into our formula {f(6)-f(4)] / (6-4)

OpenStudy (anonymous):

so -294+38? @Vocaloid

Vocaloid (vocaloid):

that's part of it the numerator is f(6) - f(4) = -294+38 and the denominator is (6-4) so divide the numerator by the denominator

OpenStudy (anonymous):

so -256/2?

Vocaloid (vocaloid):

yes

OpenStudy (anonymous):

-128?

OpenStudy (anonymous):

is that right?

OpenStudy (anonymous):

@Vocaloid

Vocaloid (vocaloid):

yup

OpenStudy (anonymous):

so the final answer is -128?

OpenStudy (anonymous):

@Vocaloid

Vocaloid (vocaloid):

yes

OpenStudy (anonymous):

AWESOME THANK YOU SO MUCH, how do i give u a medal? @Vocaloid

Vocaloid (vocaloid):

click "best response" next to my name thanks, I'm glad I could be of service

OpenStudy (anonymous):

never mind got it!

OpenStudy (anonymous):

thanks so much honesttly !